OpenStudy (anonymous):

Solve. 4(x – 3) + 5x = –48

OpenStudy (anonymous):

4(x-3)+5x=-48 4x-12+5x=-48 9x-12=-48 9x=-36 x=-4

OpenStudy (anonymous):

Solve. 7 – 3x = 3(3 – x) – 2

OpenStudy (anonymous):

ill give u a metal if u reply to this

OpenStudy (anonymous):

x=0